# TRG-Net: An Interpretable and Controllable Rain Generator

### v1

@@ 5.3.3 OOD Rain Augmentation
Rain sparsity augmentation. We select the 50mm/hr rain data from the training set of Cityscapes [30] as our training set. The test rain datasets are collected from the test sets of Cityscapes, incorporating different rainfall rates, ranging from 5mm/hr to 100mm/hr. Firstly, the rain generators are trained on 50mm/hr training set. Then, in the training phase for deraining, we use PReNet trained on original 50mm/hr training set as the baseline and further augment the training data through in-distribution augmentation using VRG-Net and TRG-Net. While for TRG-Net, out-of-distribution sparsity rain data are augmented by controlling the sparsity factor of TRG-Net, which is indicated as ${}^{\dagger}$. The deraining performance of all competing methods on all test sets are shown in Table 8. The performance of deraining on out-of-distribution sparsity rain is improved by implementing OOD rain sparsity augmentation.
Table 9: The FID and KID of generating rain by TRG-Net under different ablation settings.
Factor learning MerNet RotTV FID KID
@@
Table 10: The deraining results of PReNet[32] trained on the pseudo-paired data generated by TRG-Net under different ablation settings.
Factor learning MerNet RotTV PSNR SSIM
$\times$ 28.55 0.874
$\times$ 29.82 0.905
$\times$ 31.94 0.946
32.33 0.949
5.4 Ablation Study
@@
We conduct various ablation settings for the proposed TRG-Net. Firstly, we consider the effectiveness of the rain factor learning. Specifically, we replace transformable convolution kernels with common convolution kernels in our model, which is then degenerated to a CSC-based rain model without any rain factor input. Then we also consider the influence of the merging model and the proposed rotTV regularizer. All ablation experiments are performed on Rain100L in an unpaired way. Table 9 shows the FID and KID of generating rain by TRG-Net under different ablation settings. We can observe that the factor learning, merging model and rotTV all have effects on the rain generation performance. In particular, Table 10 further shows the deraining performance of PReNet trained on the pseudo pairs generated by these different settings. It can be observed that all the aforementioned settings make a positive contribution to the deraining performance of the entire TRG-Net model, and the completeness of them achieves the best deraining performance.
6 Conclusion
@@
In this paper, we have constructed a transformable rainy image generation network (TRG-Net), which intrinsically encodes the rain factors, including shape, orientation, length, width and sparsity, into the design of network and can properly extract these fundamental rain factors underlying rainy images purely from data. To the best our knowledge, TRG-Net should be the first rain generator that can not only elaborately design fundamental elements to simulate expected rains, like conventional artificial rendering methods, but also finely adapt to complicated and diverse practical rain patterns, like recent deep learning methods. A transformable convolution framework has been proposed to alleviate the difficulty in embedding controllable and learnable rain factors into deep networks. We have further presented a rotatable TV regularizer for rain generation, which is able to adaptively adjust the orientation for calculating variations and adopt higher penalty along the rain streak orientation. Comprehensive experiments on synthetic and real datasets have validated the superiority of the proposed TRG-Net beyond current SOTA rain generation methods in both unpaired rain generation and paired rain augmentation tasks.
Although a reasonable rainy image generator has been proposed, we find empirically that the discriminator can have an impact on the performance of rain generation. Therefore, the design of the discriminator is also critical for stabilising the training of GAN. Besides, the proposed transformable convolution framework and the rotatable TV regularizer could have potential application value for extensive tasks, such as the design of the deraining network. We will try these designs in our future research.
Appendix A More Details on The Proposed TRG-Net
Figure 12: The overall framework of the proposed TRG-Net for rain layer generation, which mainly consists of three rain kernel models and three rain map models.
Our experiments empirically show that a single layer convolution of the rain kernels with the rain maps is usually not sufficient to generate the diverse, complex and long rain streaks, therefore we actually generate three rain kernels and convolute them with the rain maps to increase the receptive field of rain kernel. The overall architecture of the proposed rain layer generation network is shown in Fig. 12. Specifically, the rain kernels in three convolutions share the same shape, orientation, length and width factors. For $i=1,2,3$,
$\mathcal{C}_{i}=\mbox{KerNet}_{\bm{w}^{c_{i}}}(e_{\theta},e_{l},e_{w},e_{\alpha}),$ (31)
where $\mbox{KerNet}(\cdot)$ is defined in the main text Eq. (15). $\bm{w}^{c_{i}}=\left\{\bm{w}^{d_{i}},\bm{w}^{\theta},\bm{w}^{l},\bm{w}^{w},\bm{w}^{\alpha}\right\}$ represent all the to-be-learnt parameters and $\{e_{\theta},e_{l},e_{w},e_{\alpha}\}\sim N(0,1)$. We can see that the parameters of rain factors ($\bm{w}^{\theta},\bm{w}^{l},\bm{w}^{w},\bm{w}^{\alpha}$) are shared, while the parameter of rain kernel dictionary $\bm{w}^{d_{i}}$ is not shared. This ensures that the rain factors implied in rain datasets are shared with all rain kernels.
The rain maps in three convolutions are all generated by rotResNet and the sparsity factor $\tau$ is encoded into first two rain map models. From Eq. (17) in the main text,
$\mathcal{M}_{1}=\mbox{ReLU}(\mbox{rotResNet}^{1}_{\bm{w}^{r_{1}}}(\mathcal{Z},\theta)-\tau),$ (32)
[…]
$\mathcal{M}_{2}\!=\!\mbox{ReLU}(\mbox{rotResNet}^{2}_{\bm{w}^{r_{2}}}(\mathcal{C}_{1}\!\otimes\!\mathcal{M}_{1},\theta)\!-\!\tau),$ (33)
where $\theta,\tau$ are shared with Eq. (32). Similarly, the rain map model takes the convolution output of the rain kernel $\mathcal{C}_{2}$ and the rain map $\mathcal{M}_{2}$ as input and outputs the final rain map. That is
$\mathcal{M}_{3}=\mbox{rotResNet}^{3}_{\bm{w}^{r_{3}}}(\mathcal{C}_{2}\otimes\mathcal{M}_{2},\theta),$ (34)
where $\theta$ is the orientation degree, which is shared with Eqs. (32) and (33).
Based on the proposed rain model (10) in the main text, the rain layer is generated by
$\mathcal{R}(\tau,\theta,s_{l},s_{w})=\mathcal{C}_{3}(\theta,s_{l},s_{w})\otimes\mathcal{M}_{3}(\tau,\theta),$ (35)
where $\mathcal{C}_{3}(\theta,s_{l},s_{w})$ is the third rain kerenl generated by the rain kernel model, and $\mathcal{M}_{3}(\tau)$ is the third rain map.
It should be noted that the rain kernel dictionary has an evident physical meaning, which is different from other convolution filters in deep networks. Therefore, we specifically design a initialization manner for it according to the prior knowledge of rains. Formally, the initialized rain kernel dictionary is sampled from
Figure 13: Some initialized rain kernel dictionaries to better understand the design of Eq. (36).
$\tilde{\phi}(x)=A\cdot\exp\left(-25\sigma\left|\left|\mbox{ReLu}(\left|x\right|-w)\right|\right|_{2}^{2}\right)$ (36)
where $\left|x\right|=\begin{bmatrix}\left|x_{1}\right|\\ \left|x_{2}\right|\end{bmatrix}$, $w=\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$, $w_{1}\sim N(0.001,0.01)$, $w_{2}\sim N(0.5,0.2)$, $\sigma\sim N(0.1,2)$, $A\sim N(0.6,0.3)$. Fig. 13 shows some initialized rain kernel dictionaries for visual understanding of this design. Then, the initialization values of the parameter $\bm{w}^{d}$ in Eq. (12) of the main text can be is easily obtained by adopting discrete Fourier transform on the initialized dictionary.
Data Availability Statement
The data supporting the results of this study are available from the authors on request.
Statements and Declarations
The authors declare that they have no known competing financial interests or personal relationships that could appear to have influenced the work reported in this paper.

### v2

@@ V-C3 OOD Rain Augmentation
The experimental results. The experimental results are shown in Table VI, where “Baseline” indicates the deraining result of PReNet trained on original training set. We also list the deraining results of two model-based methods, DSC [48] and JCAS [49]. We can see that Baseline and VRG-Net achieve a relatively poor performance when the orientation of the rain distribution in the test set differs from the training set. Comparatively, the proposed methods not only perform better on test data with orientation range in $[-30^{\circ},30^{\circ}]$, but also have a well performance on rain orientation range in ±$\pm[30^{\circ},60^{\circ}]$. Specifically, the augmented training on TRG-Net${}_{[-60^{\circ},60^{\circ}]}$ achieves a significant performance improvement in ±$\pm[30^{\circ},60^{\circ}]$ case.
TABLE VII: The FID and KID of generating rain by TRG-Net under different ablation settings.
[…]
@@
V-D Ablation Study
@@
We conduct various ablation settings for the proposed TRG-Net. Firstly, we consider the effectiveness of the rain factor learning. Specifically, we replace transformable convolution kernels with common convolution kernels in our model, which is then degenerated to a CSC-based rain model without any rain factor input. Then we also consider the influence of the merging model and the proposed rotTV regularizer [^13] For more ablation studies, please refer to the supplementary material.. All ablation experiments are performed on Rain100L in an unpaired way. Table VII shows the FID and KID of generating rain by TRG-Net under different ablation settings. It can be observed that all the aforementioned settings make a positive contribution to the deraining performance of the entire TRG-Net model, and the completeness of them achieves the best deraining performance.
VI Conclusion and Future Work
@@
In this paper, we have constructed a transformable rainy image generation network (TRG-Net), which intrinsically encodes the rain factors, including shape, orientation, length, width and sparsity, into the design of network and can properly extract these fundamental rain factors underlying rainy images purely from data without the need for rain factor labels. Its significance lies in that TRG-Net can not only elaborately design fundamental elements to simulate expected rains, like conventional artificial rendering methods, but also finely adapt to complicated and diverse practical rain patterns, like recent deep learning methods. A transformable convolution framework has been proposed to alleviate the difficulty in embedding controllable and learnable rain factors into deep networks. We have further presented a rotatable TV regularizer for rain generation, which is able to adaptively adjust the orientation for calculating variations and adopt higher penalty along the rain streak orientation. Comprehensive experiments on synthetic and real datasets have validated the superiority of the proposed TRG-Net beyond current SOTA rain generation methods in both unpaired rain generation and paired rain augmentation tasks.
While the proposed TRG-Net has encoded essential rain factors (e.g., orientation, width, length and density) into the design of the rain generation network in a controllable and interpretable way, other important factors, such as depth and photometry, are simply embedded by the ResNet-based merging network. Therefore, the design of the merging network and its integration with modules proposed in previous works should be further explored, especially when applying the proposed method in practical scenarios. Besides, the proposed transformable convolution framework and the rotatable TV regularizer could have potential application value for extensive tasks, such as the design of the deraining network. We will try these designs in our future research.